A cuvette of length l is filled with a solution of aspirin with concentration c. You are recording the transmission of light of a certain wavelength and record a transmission of T=90%. A= -log(%T / 100%), A=\varepsilon*b*c. Show your work to receive credit.

a) What is the absorbance of the solution?

b) The concentration of aspirin is increased by a factor of 2. What is the absorbance of the cell? What is the transmission of the cell?

c) The concentration of aspirin is decreased by a factor of 2. What is the absorbance of the cell? What is the transmission of the cell?
Answer #1

A = εbc

ε = molar absorptivity = liters/mole cm

b = cell path length = cm

c = concentration = moles/liter

A= -log(%T / 100%) or A = log10(100/T) = log10(100) - log10 (T) = 2 - log10 (T)

a) Given T = 90% = 90/100 = 0.9, A = 2 - log10 (T) = 2 - log10 (0.9) = 1.95

absorbance of the solution = 1.95

b) If the solution concentration increases by a factor of 2 then the absorbance will also increase by a factor of 2.

absorbance of the cell = 1.95*2 = 3.9

transmission of the cell T = 10-A = 10-3.9= 0.01259%

c) If the solution concentration decreases by a factor of 2 then the absorbance will also decrease by a factor of 2.

absorbance of the cell = 1.95/2 = 0.975

transmission of the cell T = 10-A = 10-0.975= 10.59%